# Efficient Deep Learning with Decorrelated Backpropagation

### v4

@@ Abstract
The backpropagation algorithm remains the dominant and most successful method for training deep neural networks (DNNs). At the same time, training DNNs at scale comes at a significant computational cost and therefore a high carbon footprint. Converging evidence suggests that input decorrelation may speed up deep learning. However, to date, this has not yet translated into substantial improvements in training efficiency in large-scale DNNs. This is mainly caused by the challenge of enforcing fast and stable network-wide decorrelation. Here, we show for the first time that much more efficient training of deep convolutional neural networks is feasible by embracing decorrelated backpropagation as a mechanism for learning. To achieve this goal we made use of a novel algorithm which induces network-wide input decorrelation using minimal computational overhead. By combining this algorithm with careful optimizations, we achieve a more than two-fold speed-up and higher test accuracy compared to backpropagation when training several deep networks up to a 50-layer ResNet model. This demonstrates that decorrelation provides exciting prospects for efficient deep learning at scale.
@@ 4 Discussion
Increasing the efficiency of deep learning is of key importance if we are to reduce the carbon footprint of AI. In this paper, using four deep neural network architectures trained on the Imagenet dataset, we have shown that decorrelated backpropagation, which decorrelates the inputs to each network layer, provides a viable path towards more efficient deep learning. Results show that, by replacing BP with DBP, we achieve higher test accuracy using up to half the number of training epochs, resulting in reduced carbon emission.

### v5

@@ Abstract
The backpropagation algorithm remains the dominant and most successful method for training deep neural networks (DNNs). At the same time, training DNNs at scale comes at a significant computational cost and therefore a high carbon footprint. Converging evidence suggests that input decorrelation may speed up deep learning. However, to date, this has not yet translated into substantial improvements in training efficiency in large-scale DNNs. This is mainly caused by the challenge of enforcing fast and stable network-wide decorrelation. Here, we show for the first time that much more efficient training of deep convolutional neural networks is feasible by embracing decorrelated backpropagation as a mechanism for learning. To achieve this goal we made use of a novel algorithm which induces network-wide input decorrelation using minimal computational overhead. By combining this algorithm with careful optimizations, we achieve a more than two-fold speed-up and higher test accuracy compared to backpropagation when training several deep residual networks. This demonstrates that decorrelation provides exciting prospects for efficient deep learning at scale.
@@ 3.5 Impact of whitening
3.6 Robustness to data regimes
The results in Sections 3.2 and 3.3 intentionally show performance in a setup without data augmentation or other methods to aid performance, as the aim was to make a clean comparison of the convergence properties of the algorithms. In this section, we will test the robustness of our findings by exposing both BP and DBP to two additional data regimes: data augmentation consisting of random cropping and horizontal flipping as described in He et al. (2016), as well as downsampling the training set to 10%, stratified over classes.
Figure 8: Test performance of BP and DBP ($\kappa=0.5$) on ResNet18 using for the regular ImageNet dataset (solid line), data augmentation (dashed line) and data downsampled to 10% (dotted line).
Results in Figure 8 indicate that DBP’s outperformance of BP is robust in all three data regimes. Data augmentation improved test performance over baseline for both algorithms, which now reach 65% top-1 test accuracy, close to published benchmarks He et al. (2016). Data downsampling obviously hurt test performance for both algorithms, but DBP’s higher maximum test accuracy and faster convergence remain intact.
@@ 4 Discussion
Increasing the efficiency of deep learning is of key importance if we are to reduce the carbon footprint of AI. In this paper, using four deep neural network architectures trained on the Imagenet dataset, we have shown that decorrelated backpropagation, which decorrelates the inputs to each network layer, provides a viable path towards more efficient deep learning. Results show that, by replacing BP with DBP, we achieve higher test accuracy using up to half the number of training epochs, resulting in reduced carbon emission. Results appear to be robust under both data augmentation and data downsampling.